OpenStudy (anonymous):

If E and F are events with P(E) = 0.4 and P(F) = 0.3 and P(E n F) = 0.2, find the probability of E given F. Also find P(F I E)

OpenStudy (zarkon):

\[P(E|F)=\frac{P(E\cap F)}{P(F)}\]

OpenStudy (zarkon):

\[P(F|E)=\frac{P(F\cap E)}{P(E)}\]

myininaya (myininaya):

just plug into that second formula zarkon has :) you can do it

OpenStudy (anonymous):

so P(F I E) = .5

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

and how would i find P( E I F)

myininaya (myininaya):

use the very first formula he has for you

OpenStudy (anonymous):

so .67

OpenStudy (anonymous):

?

OpenStudy (zarkon):

\[\frac{2}{3}\]

OpenStudy (zarkon):

\[.\bar{6}\]

OpenStudy (anonymous):

yep thats what i got!!!! thank you!!!!!!!!!!!

OpenStudy (anonymous):

.6ˉ